### v2

@@ 3.2 Examples
Li et al. (2023a) provide a series of concrete examples satisfying the $\ell$-smooth condition. Since our formulation generalizes theirs, these example functions also satisfy $\ell*$-smoothness. Nonetheless, we provide additional examples to illustrate the potential emergence of dimension-dependent factors, highlighting the advantages of the algorithmic adaptivity brought by mirror descent. The comprehensive discussions and omitted details can be found in Appendix B.
@@
$f_{1}(\mathbf{x}):=\exp\left(\mathbf{w}^{\top}\mathbf{x}+b\right)\quad\text{and}\quad f_{2}(\mathbf{x})=\log\left(1+\exp(-\mathbf{w}^{\top}\mathbf{x})\right)$ (4)
@@
Proposition 4.
  1. (i)
    $f_{1},f_{2}\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with ^$\widehat{\ell}(\alpha)=\|\mathbf{w}\|_{2}\,\alpha$.
  2. (ii)
    $f_{1},f_{2}\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with ~$\widetilde{\ell}(\alpha)=\|\mathbf{w}\|_{\infty}\,\alpha$.
  3. (iii)
[…]
4 shows that our $\ell*$-smoothness can be significantly tighter than $\ell$-smoothness of Li et al. (2023a), with a worst-case dimensional factor of $1/\sqrt{n}$. In Sections B.2 and B.3, we further exhibit examples where ~$\widetilde{\ell}(\alpha)/\widehat{\ell}(\alpha)=O(1/n)$ and $O(n^{-0.4})$, respectively, demonstrating even stronger gains. Since both Li et al. (2023a) and our Theorem 8 yield convergence rates proportional to the link function $\ell$, mirror descent under $\ell*$-smoothness achieves improved rates by these dimension-dependent factors (see Section B.4 for detailed discussions).
@@
The examples and the justifications in 4 are not confined to unbounded domains like $\mathbb{R}^{n}$. They remain valid under Assumption 2, as elaborated in Section B.1.3. Moreover, we underline that the constant-link function example in Section B.2 as well as the empirical example in Section B.3 are defined on bounded domains well-modeled by Assumption 2.
4 Mirror Descent Meets Generalized Smoothness
This section formally presents the convergence results for a series of first-order optimization algorithms based on mirror descent, whose analyses can be found in Section 6.
@@ 4.1 Preliminaries
Definition 5.
[…]
Definition 6.
Define the Bregman function $B:\mathcal{X}\times\mathcal{X}^{o}\mapsto\mathbb{R}_{+}$ associated with the distance-generating function $\psi$ as
$B(\mathbf{x},\mathbf{x}^{o})=\psi(\mathbf{x})-\psi(\mathbf{x}^{o})-\langle\nabla\psi(\mathbf{x}^{o}),\mathbf{x}-\mathbf{x}^{o}\rangle,\quad\forall(\mathbf{x},\mathbf{x}^{o})\in\mathcal{X}\times\mathcal{X}^{o}.$ (5)
[…]
Assumption 2
The domain $\mathcal{X}$ is bounded in the sense that
$\max_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})-\min_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})\leq D^{2},\quad D\in\mathbb{R}_{++}.$ (6)
Assumption 3
The objective function satisfies $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$, where $\ell$ is sub-quadratic, i.e., $\lim_{\alpha\to\infty}\allowbreak\alpha^{2}/\ell(\alpha)=\infty$.
For convenience, we also define the following prox-mapping function (Nemirovski, 2004):
$\mathcal{P}_{\mathbf{y}}(\mathbf{g})=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}\left\{\langle\mathbf{g},\mathbf{x}\rangle+B(\mathbf{x},\mathbf{y})\right\},\quad\forall\mathbf{y}\in\mathcal{X},\mathbf{g}\in\mathcal{E}^{*},$ (7)
@@ 4.2 Main Idea and Technical Challenges
As discussed previously, bounding the gradients is a prerequisite for Lemma 3, which enables us to reduce the $\ell*$-smoothness analysis to the conventional analysis of $L$-smooth functions. In the Euclidean case, Li et al. (2023a) bound the gradients along the optimization trajectory by tracking the gradients directly. More specifically, they manage to show that for gradient descent, the $\ell_{2}$-norm of the gradients decrease monotonically provided that the learning rates are sufficiently small. However, this vital discovery builds upon the geometry of the Euclidean setting, i.e., $\langle\mathbf{x},\mathbf{x}\rangle=\left\|\mathbf{x}\right\|_{2}^{2}$, which is invalid in the non-Euclidean setting.
Technical challenges
The underlying problem is that for mirror descent, the dual norms of the gradients do not necessarily form a monotonic decreasing sequence. The difficulty stems from the nature of mirror descent, where the descent step is performed in the dual space, making the explicit analysis of the dual norms very difficult and rarely addressed. Even under $L$-smoothness, tracking the gradients directly for mirror descent remains challenging, as discussed in Zhang and He (2018); Lan (2020); Huang et al. (2021) and references therein, unless the problem exhibits special structures (Zhou et al., 2017) or extra assumptions are exerted (Lei and Zhou, 2017).
Main idea
To tackle this problem, we introduce the following generalized version of the reversed Polyak-Łojasiewicz inequality (Polyak, 1963; Lojasiewicz, 1963), aka the self-bounding property of smooth functions (Srebro et al., 2010; Zhang et al., 2017; Zhang and Zhou, 2019).
Lemma 7.
@@
Lemma 7 measures the growth of gradient’s dual norm by the suboptimality gap $f(\mathbf{x})-f^{*}$. By Assumption 3, if we further assume $\ell(\alpha)=(\alpha/2)^{\beta},\beta\in[0,2)$, then $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq 2[f(\mathbf{x})-f^{*}]^{1/(2-\beta)}$ is immediately implied. Hence, Lemma 7 suggests that $\left\|\nabla f(\mathbf{x})\right\|_{*}$ is effectively under control if $\ell(\cdot)$ does not grow too rapidly. In this way, the gradient can be bounded by the suboptimality gap of the objective function, which is generally much easier in the context of mirror descent. We rigorously show that the suboptimality gap at an arbitrary iterate is bounded by that of the starting point (see Lemma 18), which further suggests the boundness of the gradients. Once the gradients have an upper bound, we can construct an effective smoothness parameter $L$ that serves as an optimistic estimation of the local curvatures along the trajectory. Formally, we define
$G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2\ell(2\alpha)(f(\mathbf{x}_{0})-f^{*})\right\},\text{ and }L:=\ell(2G),$ (8)
where $\mathbf{x}_{0}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})$ is selected as the starting point for all optimization algorithms considered in this paper, whose trajectory provably satisfies $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$ (will be stated rigorously in the theorems to come) and consequently, $\ell(2\left\|\nabla f(\mathbf{x}_{t})\right\|_{*})\leq L$. We also emphasize that both $G$ and $L$ are absolute constants, i.e., $G,L<\infty$ (Lemma 16), which is essential to the validity of our theoretical derivations.
@@ 4.3 Mirror Descent and Accelerated Mirror Descent
where $\eta$ represents the learning rates, and the prox-mapping function $\mathcal{P}_{\cdot}(\cdot)$ is defined in (7). Our convergence analysis relies on the interesting fact that the suboptimality gap along the trajectory is a non-increasing sequence: $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{t-1})-f^{*}\leq\cdots\leq f(\mathbf{x}_{0})-f^{*}$, provided that learning rates are chosen appropriately. Consequently, the gradients along the trajectory are bounded by Lemma 7, and so are the local Lipschitz constants. The formal statements are illustrated in the following theorem.
Theorem 8.
Under Assumptions 1, 2 and 3, if $0<\eta\leq 1/L$, then the gradients along the trajectory generated by (9) satisfy $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G,\forall t\in\mathbb{N}$, and the average-iterate convergence rate is given by
$f(\overline{\mathbf{x}}_{T})-f^{*}\leq\frac{D^{2}}{\eta T},$ (10)
where $\overline{\mathbf{x}}_{T}=\sum_{t=1}^{T}\mathbf{x}_{t}/T$. Moreover, the last-iterate convergence rate is given by
$f(\mathbf{x}_{T})-f^{*}\leq\frac{D^{2}}{\eta T}.$ (11)
Our convergence result recovers the classical $O(1/T)$ rate for deterministic convex optimization (Bubeck et al., 2015; Lan, 2020), for both the average-iterate and the last-iterate.
@@
We present the following theoretical guarantee, which achieves the optimal $O(1/T^{2})$ convergence rate of first-order methods (Nesterov et al., 2018).
Theorem 9.
Under Assumptions 1, 2 and 3, define [^3] Due to technical reasons, $L=\ell(4G)$ needs to be slightly larger than the one used in mirror descent. This adjustment affects only accelerated mirror descent; other algorithms still use $L=\ell(2G)$.
$L:=\ell(4G),\tau:=\left\lceil\frac{4\sqrt{2}DL}{G}\right\rceil,\eta:=\min\left\{1,\frac{3(\tau-1)}{2(\tau-3)(\tau-2)}\right\},\eta_{t}:=\frac{t\eta}{2L},\alpha_{t}:=\frac{2}{t+1}.$ (13)
@@
$f(\mathbf{x}_{T})-f^{*}\leq\frac{D^{2}L}{\eta T(T+1)},$ (14)
@@
The general idea is similar to mirror descent, that is, bounding the gradients along the trajectory via controlling the suboptimality gaps. However, as (12) maintains three sequences, the analysis becomes significantly more challenging compared to the single sequence $\left\{\mathbf{x}_{t}\right\}_{t\in\mathbb{N}}$ in (9). While $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\lesssim G$ follows directly from the convergence in (14), it cannot be directly extended to the sequence $\{\mathbf{y}_{t}\}_{t\in\mathbb{N}}$, whose gradients also need to be bounded. To address this issue, we focus on the key quantity $e_{t}:=\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|$, which measures the distance between the two sequences. If $e_{t}\lesssim G/L$, we can readily deduce $\left\|\nabla f(\mathbf{y}_{t})\right\|_{*}\lesssim G$ by leveraging the local smoothness property in Definition 1. To achieve this goal, we develop a “time partition” technique illustrated in Lemma 21. Specifically, when $t$ is below the threshold $\tau$, we bound $e_{t}$ by a contraction mapping, which effectively limits its growth for small $t$. Conversely, when $t$ exceeds $\tau$, $e_{t}$ provably decays hyperbolically. Combining the two cases, we can derive that $e_{t}\lesssim G/L$, which suggests the boundness of the gradients.
Comparison with Li et al. (2023a, Theorem 4.4)
For $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|_{2})$Li et al. (2023a) study a variant of Nesterov’s accelerated gradient algorithm (NAG) and also recover the optimal bound. They introduce an auxiliary sequence into the original NAG, which aims to aggressively stabilize the optimization trajectory. The stabilization mainly refers to $\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|$ in their language, sharing a similar spirit as $e_{t}$ in our paper. To this end, they have to use a smaller step size ($\eta\simeq 1/L^{2}$) and impose a more complex analysis than the original NAG. Unlike their framework, we concentrate on a more stable acceleration scheme in (12), without manually tuning down the base learning rate $\eta$ or bringing in any additional algorithmic components. Moreover, our proof is much simpler and more intuitive, as reflected in our constructive ways of analyzing the stability term $e_{t}$ by partitioning the timeline. It is also worth mentioning that we do not need the fine-grained characterization on the modulus of $\ell$, while Li et al. (2023a) further assume $\ell(u)=O(u^{\alpha}),\alpha\in(0,2]$.

### v3

@@ 2.2 Examples
Li et al. (2023a) provide a series of concrete examples satisfying the $\ell$-smooth condition. Since our formulation generalizes theirs, these example functions also satisfy $\ell*$-smoothness. Nonetheless, we provide additional examples to illustrate the potential emergence of dimension-dependent factors, highlighting the advantages of the algorithmic adaptivity brought by mirror descent. The comprehensive discussions and omitted details can be found in Appendix C.
@@
$f_{1}(\mathbf{x}):=\exp\left(\mathbf{w}^{\top}\mathbf{x}+b\right),f_{2}(\mathbf{x})=\log\left(1+\exp(-\mathbf{w}^{\top}\mathbf{x})\right)$
@@
Proposition 2.8.
(i) $f_{1},f_{2}\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\alpha)=\|\mathbf{w}\|_{2}\alpha$. (ii) $f_{1},f_{2}\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\alpha)=\|\mathbf{w}\|_{\infty}\alpha$. (iii) The ratio of link functions satisfies $1/\sqrt{n}\leq\widetilde{\ell}(\alpha)/\widehat{\ell}(\alpha)\leq 1$.
Proposition 2.8 shows that our $\ell*$-smoothness can be significantly tighter than $\ell$-smoothness of Li et al. (2023a), with a worst-case dimensional factor of $1/\sqrt{n}$. In Sections C.2 and C.3, we further exhibit examples where $\widetilde{\ell}(\alpha)/\widehat{\ell}(\alpha)=O(1/n)$ and $O(n^{-0.4})$, respectively, demonstrating even stronger gains. Since both Li et al. (2023a) and our Theorem 3.5 yield convergence rates proportional to the link function $\ell$, mirror descent under $\ell*$-smoothness achieves improved rates by these dimension-dependent factors (see Section C.4 for detailed discussions).
@@
The examples and the justifications in Proposition 2.8 are not confined to unbounded domains like $\mathbb{R}^{n}$. They remain valid under Assumption J.2, as elaborated in Section C.1.3. Moreover, we underline that the constant-link function example in Section C.2 as well as the empirical example in Section C.3 are defined on bounded domains well-modeled by Assumption J.2.
2.3 Empirical Demonstration
Consider the layer-wise non-Euclidean smoothness model proposed in Riabinin et al. (2025, Assumption 1):
$\frac{\|\nabla_{i}f(X)-\nabla_{i}f(Y)\|_{(i)\star}}{\|X_{i}-Y_{i}\|_{(i)}}\leq L^{0}_{i}+L^{1}_{i}\|\nabla_{i}f(X)\|_{(i)\star},$ (3)
where $i\in[K]$ denotes the $i$-th layer of a deep neural network. Notice that our Definitions 2.1 and 2.4 actually cover the above assumption since setting $\left\|\cdot\right\|$ in (2) to $\left\|\cdot\right\|_{(i)}$ directly generates (3). Motivated by this, we experiment on nanoGPT pretraining to demonstrate the address the $\ell*$-smoothness, with results shown in Figure 1. A strong agreement between real and approximated $L_{i}$ values suggests that our $\ell*$-smooth assumption aligns well with the real optimization trajectory of LLMs. Further experimental details can be found in Appendix B.
Figure 1: Trajectory of NanoGPT training on FineWeb dataset
3 Mirror Descent Meets $\ell*$-smoothness
This section formally presents the convergence results for a series of first-order optimization algorithms based on mirror descent, whose analyses can be found in Appendices E, F, G and H.
@@ 3.1 Preliminaries
Definition 3.1.
We call a continuous function $\psi:\mathcal{X}\mapsto\mathbb{R}_{+}$ a distance-generating function with modulus $\alpha$ w.r.t. $\left\|\cdot\right\|$, if (i) the set $\mathcal{X}^{o}=\{\mathbf{x}\in\mathcal{X}|\partial\psi(\mathbf{x})\neq\emptyset\}$ is convex; (ii) $\psi$ is continuously differentiable and $\alpha$-strongly convex w.r.t. $\left\|\cdot\right\|$, i.e., $\langle\nabla\psi(\mathbf{x}_{1})-\nabla\psi(\mathbf{x}_{2}),\mathbf{x}_{1}-\mathbf{x}_{2}\rangle\geq\alpha\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|^{2},\ \forall\mathbf{x}_{1},\mathbf{x}_{2}\allowbreak\in\mathcal{X}^{o}$.
Definition 3.2.
Define the Bregman function $B:\mathcal{X}\times\mathcal{X}^{o}\mapsto\mathbb{R}_{+}$ associated with the distance-generating function $\psi$ as $B(\mathbf{x},\mathbf{x}^{o})=\psi(\mathbf{x})-\psi(\mathbf{x}^{o})-\langle\nabla\psi(\mathbf{x}^{o}),\mathbf{x}-\mathbf{x}^{o}\rangle$.
We equip domain $\mathcal{X}$ with a distance-generating function $\psi(\cdot)$, which is, WLOG, $1$-strongly convex w.r.t. a certain norm $\left\|\cdot\right\|$ endowed on $\mathcal{E}$. We also define the corresponding Bregman divergence according to Definition 3.2. Then, we introduce the following assumptions on the objective function, which controls the growth rate of $\ell(\cdot)$, as also required in Li et al. (2023a).
Assumption 3.3.
The objective satisfies $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$, where $\ell$ is sub-quadratic, i.e., $\lim_{\alpha\to\infty}\allowbreak\alpha^{2}/\ell(\alpha)=\infty$.
For convenience, we also define the following prox-mapping function (Nemirovski, 2004). For any $\mathbf{g}\in\mathcal{E},\mathbf{y}\in\mathcal{X}$,
$\mathcal{P}_{\mathbf{y}}(\mathbf{g}):=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}\left\{\langle\mathbf{g},\mathbf{x}\rangle+B(\mathbf{x},\mathbf{y})\right\},$ (4)
@@ 3.2 Main Idea and Technical Challenges
As discussed previously, bounding the gradients is a prerequisite for Lemma 2.7, which enables us to reduce the $\ell*$-smoothness analysis to the conventional analysis of $L$-smooth functions. In the Euclidean case, Li et al. (2023a) bound the gradients along the optimization trajectory by tracking the gradients directly. More specifically, they manage to show that for gradient descent, the $\ell_{2}$-norm of the gradients decrease monotonically provided that the learning rates are sufficiently small. However, this vital discovery builds upon the geometry of the Euclidean setting, i.e., $\langle\mathbf{x},\mathbf{x}\rangle=\left\|\mathbf{x}\right\|_{2}^{2}$, which is invalid in the non-Euclidean setting.
Technical challenges   The underlying problem is that for mirror descent, the dual norms of the gradients do not necessarily form a monotonic decreasing sequence. The difficulty stems from the nature of mirror descent, where the descent step is performed in the dual space, making the explicit analysis of the dual norms very difficult and rarely addressed. Even under $L$-smoothness, tracking the gradients directly for mirror descent remains challenging, as discussed in Zhang and He (2018); Lan (2020); Huang et al. (2021) and references therein, unless the problem exhibits special structures (Zhou et al., 2017) or extra assumptions are exerted (Lei and Zhou, 2017).
Main idea   To tackle this problem, we introduce the following generalized version of the reversed Polyak-Łojasiewicz inequality (Polyak, 1963; Lojasiewicz, 1963), aka the self-bounding property of smooth functions (Srebro et al., 2010; Zhang et al., 2017; Zhang and Zhou, 2019).
Lemma 3.4.
@@
Lemma 3.4 measures the growth of gradient’s dual norm by the suboptimality gap $f(\mathbf{x})-f^{*}$. By Assumption 3.3, if we further assume $\ell(\alpha)=(\alpha/2)^{\beta},\beta\in[0,2)$, then $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq 2[f(\mathbf{x})-f^{*}]^{1/(2-\beta)}$ is immediately implied. Hence, Lemma 3.4 suggests that $\left\|\nabla f(\mathbf{x})\right\|_{*}$ is effectively under control if $\ell(\cdot)$ does not grow too rapidly. In this way, the gradient can be bounded by the suboptimality gap of the objective function, which is generally much easier in the context of mirror descent. We rigorously show that the suboptimality gap at an arbitrary iterate is bounded by that of the starting point (see Lemma E.2), which further suggests the boundness of the gradients. Once the gradients have an upper bound, we can construct an effective smoothness parameter $L$ that serves as an optimistic estimation of the local curvatures along the trajectory. Formally, we define
$G=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2\ell(2\alpha)(f(\mathbf{x}_{0})-f^{*})\right\},$ (5)
$\text{ and }L=\ell(2G),$
where $\mathbf{x}_{0}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})$ is selected as the starting point for all optimization algorithms considered in this paper, whose trajectory provably satisfies $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G$ (will be stated rigorously in the theorems to come) and consequently, $\ell(2\left\|\nabla f(\mathbf{x}_{t})\right\|_{*})\leq L$. We also emphasize that both $G$ and $L$ are absolute constants, i.e., $G,L<\infty$ (Lemma D.3), which is essential to the validity of our theoretical derivations.
@@ 3.3 Mirror Descent and Accelerated Mirror Descent
where $\eta$ represents the learning rates, and the prox-mapping function $\mathcal{P}_{\cdot}(\cdot)$ is defined in (4). Our convergence analysis relies on the interesting fact that the suboptimality gap along the trajectory is a non-increasing sequence: $f(\mathbf{x}_{t})-f^{*}\leq f(\mathbf{x}_{t-1})-f^{*}\leq\cdots\leq f(\mathbf{x}_{0})-f^{*}$, provided that learning rates are chosen appropriately. Consequently, the gradients along the trajectory are bounded by Lemma 3.4, and so are the local Lipschitz constants. The formal statements are illustrated in the following theorem.
Theorem 3.5.
Under Assumptions 2.5 and 3.3, if $0<\eta\leq 1/L$, then the gradients along the trajectory generated by (6) satisfy $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\leq G,\forall t\in\mathbb{N}$, and the average-iterate as well as the last-iterate convergence rate is given by
$\max\left\{f(\overline{\mathbf{x}}_{T})-f^{*},f(\mathbf{x}_{T})-f^{*}\right\}\leq\frac{B(\mathbf{x}_{*},\mathbf{x}_{0})}{\eta T},$ (7)
where $\overline{\mathbf{x}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{x}_{t}$.
Our convergence result recovers the classical $O(1/T)$ rate for deterministic convex optimization (Bubeck and others, 2015; Lan, 2020), for both the average-iterate and the last-iterate.
Remark 3.6.
The fact that $G$ is a finite constant is proven after we analyze the trajectory, rather than being used as an explicit assumption to analyze the trajectory. At first glance, it will incur a “circular” analysis. However, this seemingly circularity is exactly the key to our analysis techniques, where we break rigorously in Lemmas D.3, E.2 and 3.5 by induction.
@@
We present the following theoretical guarantee, which achieves the optimal $O(1/T^{2})$ convergence rate of first-order methods (Nesterov and others, 2018).
Theorem 3.7.
Under Assumptions 2.5 and 3.3, define [^2] Due to technical reasons, $L=\ell(4G)$ needs to be slightly larger than the one used in mirror descent. This adjustment affects only accelerated mirror descent; other algorithms still use $L=\ell(2G)$.
$L=\ell(4G),\tau=\left\lceil\frac{4\sqrt{2B(\mathbf{x}_{*},\mathbf{x}_{0})}L}{G}\right\rceil-1,\alpha_{t}=\frac{2}{t+1},$
$\eta=\min\left\{1,\frac{3(\tau-1)}{2(\tau-3)(\tau-2)}\right\},\eta_{t}=\frac{t\eta}{2L}.$
@@
$f(\mathbf{x}_{T})-f^{*}\leq\frac{LB(\mathbf{x}_{*},\mathbf{x}_{0})}{\eta T(T+1)},$ (9)
@@
The general idea is similar to mirror descent, that is, bounding the gradients along the trajectory via controlling the suboptimality gaps. However, as (8) maintains three sequences, the analysis becomes significantly more challenging compared to the single sequence $\left\{\mathbf{x}_{t}\right\}_{t\in\mathbb{N}}$ in (6). While $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}\lesssim G$ follows directly from the convergence in (9), it cannot be directly extended to the sequence $\{\mathbf{y}_{t}\}_{t\in\mathbb{N}}$, whose gradients also need to be bounded. To address this issue, we focus on the key quantity $e_{t}:=\left\|\mathbf{y}_{t}-\mathbf{x}_{t-1}\right\|$, which measures the distance between the two sequences. If $e_{t}\lesssim G/L$, we can readily deduce $\left\|\nabla f(\mathbf{y}_{t})\right\|_{*}\lesssim G$ by leveraging the local smoothness property in Definition 2.4. To achieve this goal, we develop a “time partition” technique illustrated in Lemma F.3. Specifically, when $t$ is below the threshold $\tau$, we bound $e_{t}$ by a contraction mapping, which effectively limits its growth for small $t$. Conversely, when $t$ exceeds $\tau$, $e_{t}$ provably decays hyperbolically. Combining the two cases, we can derive that $e_{t}\lesssim G/L$, which suggests the boundness of the gradients.
Comparison with Li et al. (2023a, Theorem 4.4)   For $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|_{2})$Li et al. (2023a) study a variant of Nesterov’s accelerated gradient algorithm (NAG) and also recover the optimal bound. They introduce an auxiliary sequence into the original NAG, which aims to aggressively stabilize the optimization trajectory. The stabilization mainly refers to $\left\|\mathbf{y}_{t}-\mathbf{x}_{t}\right\|$ in their language, sharing a similar spirit as $e_{t}$ in our paper. To this end, they have to use a smaller step size ($\eta\simeq 1/L^{2}$) and impose a more complex analysis than the original NAG. Unlike their framework, we concentrate on a more stable acceleration scheme in (8), without manually tuning down the base learning rate $\eta$ or bringing in any additional algorithmic components. Moreover, our proof is much simpler and more intuitive, as reflected in our constructive ways of analyzing the stability term $e_{t}$ by partitioning the timeline. It is also worth mentioning that we do not need the fine-grained characterization on the modulus of $\ell$, while Li et al. (2023a) further assume $\ell(u)=O(u^{\alpha}),\alpha\in(0,2]$.